OpenStudy (anonymous):

perform the indicated operation and if possible simplify . final ans as a fraction in lowest terms 25/48 / 15/22

OpenStudy (anonymous):

\[\frac{25}{48}\div\frac{15}{22}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

than filp and time \[\frac{25}{48}\times\frac{22}{15}\]can you solve now

OpenStudy (anonymous):

i can only posible do it if the number on the botton are the same i get confuse when they ar different

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

\[\frac{55\times22}{48\times15}=\frac{550}{720}=\frac{55}{72}\]

OpenStudy (anonymous):

is your answer

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

uwc

OpenStudy (anonymous):

would u help me with this one is almost the same but addtion